# The Case of Qaraaoun Reservoir, Lebanon

### v1

@@ 5 Conclusion
This study proposes an innovative approach to monitor water volume of the Qaraaoun Reservoir in real-time through the utilization of open-source satellite imagery, along with the use of a well-documented solid pipeline for water extent detection and advanced machine learning algorithms. Through thorough examination, highly accurate segmentation and reliable volume estimations were achieved. The results demonstrated how our methods contributed to building a cost-effective, sensor-free solution with minimal error and resilience against environmental challenges. Despite its promising outcomes, this solution presents certain challenges and limitations that come with the reliance on satellite data, such as atmospheric interference, which has a direct impact on segmentation accuracy. In addition to that, the chosen water index must be well-tailored in accordance with the local geographic features and seasonal changes and mechanism of water inflow into the reservoir. However, by combining multiple indices and enhancing the inference model, we were able to pave the way for durable monitoring systems.
The Qaraaoun Reservoir is a vital source for Lebanon’s agriculture (more than 40.000 ha) and water supply, which stands to benefit greatly from this tool, especially with limited infrastructure and sensor malfunctions. This research highlights the potential of remote sensing and machine learning in overcoming local challenges and advancing water management practices for developing countries. Moving forward, further optimization of the algorithm and model, alongside more extensive real-time data integration, could contribute to the long-term sustainability and resilience of the reservoir, ensuring its ability to support Lebanon’s water needs amid changing climatic conditions. Also, the integration of water quality indices extracted from LandSat and Sentinel imagery would be a valuable addition to the proposed dashboard.

### v2

@@ 5 Conclusion
This study proposes an innovative approach for real-time monitoring of the Qaraaoun Reservoir’s water volume through the utilization of open-source satellite imagery, supported by a robust and well-documented pipeline for water extent detection and advanced machine learning algorithms. Through comprehensive analysis, highly accurate segmentation and reliable volume estimations were achieved. The results demonstrate that the proposed methodology enables a cost-effective, sensor-free solution characterized by minimal error and strong resilience to environmental challenges.
Despite its promising outcomes, the approach presents certain limitations inherent to the reliance on satellite data, such as atmospheric interference, which directly affects segmentation accuracy. Furthermore, the selected water index must be carefully adapted to local geographic conditions, seasonal variations, and the specific hydrological mechanisms governing inflows into the reservoir. Nevertheless, by combining multiple indices and enhancing the inference model, this study has paved the way toward durable and scalable monitoring systems.
The Qaraaoun Reservoir represents a vital resource for Lebanon’s agriculture (supporting more than 40,000 ha) and domestic water supply, both of which stand to benefit significantly from this proposed tool—particularly given the country’s limited monitoring infrastructure and frequent sensor malfunctions. This research underscores the potential of remote sensing and machine learning to address local challenges and to advance sustainable water management practices in developing regions. Moving forward, further optimization of the algorithm and inference model, together with the integration of more extensive real-time datasets, could enhance the system’s long-term sustainability and operational resilience. Additionally, incorporating water quality indices derived from Landsat and Sentinel imagery would represent a valuable extension to the proposed dashboard, enriching its analytical and environmental monitoring capabilities.